# One More Glance with Sharp Eyes: Rethinking Lightweight Captioning as a Practical Visual Specialist

### v1

@@ D.2 Limitations of Existing Multimodal Large Language Models
We examine the broad challenges faced by existing multimodal large language models (MLLMs), particularly in addressing the issue of visual blindness. Prior works such as Eyes Wide Shut [66] and Cambrian [65] have identified and attempted to mitigate this issue through the use of multiple vision encoders, including DINOv2 [53], SigLIPv2 [68], and CLIP [56]. However, as illustrated in Fig. 13, even models with large parameter sizes continue to exhibit difficulty in generating consistent long-form captions in complex, multi-object scenes. We also conduct experiments using two recent capable MLLMs, LLaVA-Next [36] and LLaVA-OneVision [34]. Despite incorporating the sophisticated approach which is to split an image into grids and utilize all the features from each grid, both models continue to suffer from incorrect captioning results.
@@
These findings suggest that visual blindness is a persistent limitation across model scales and architectures. In this context, we believe that our Sharp-Eyed Refinement framework—which encourages attention to key regions via initial captions and leverages multi-level features from a single vision encoder—as a meaningful and efficient step toward addressing this challenge.
D.3 Limitations of lightweight captioners
@@
While we have demonstrated the potential of small models in captioning tasks, the use of lightweight LMs unavoidably introduces some limitations. In particular, we observe that the model occasionally suffers from issues such as repetitive phrasing, reduced fluency, limited OCR capability, and a lack of general world knowledge. Examples of these cases are provided in Fig. 14. These limitations may stem from two primary factors: (i) the small number of parameters, which can restrict the model’s capacity for complex reasoning and language generation [30], and (ii) the limited scale and quality of training data, as our model was trained on approximately 500K image-caption pairs from ShareGPT-4V [11] which contains machine-generated captions. A natural direction for future work is to investigate how far the capabilities of small models can be scaled with access to larger and higher-quality training datasets. In addition to the results in Sec. D.2, we observe similar issues in larger models such as LLaVA-1.5, suggesting that these challenges remain unresolved [5, 66] and require deeper investigation.
D.4 Limitations of existing evaluation methods
@@
To ensure fair comparison across captioning models, we adopt seven evaluation metrics: BLEU@4 [55], METEOR [16], CIDEr [71], BERTScore [93], CAPTURE [18], CLAIR [8], and MLLM-as-judge [9] (i.e., GPT-based evaluation). For CLAIR and MLLM-as-judge, we randomly select 100 image samples and evaluate them using 10 different seeds to compute both mean and standard deviation. The values reported alongside the $\pm$ symbol indicate the standard deviation across seeds. Despite these efforts, we find that current evaluation methods do not always align well with human judgment. For instance, certain models perform better under specific metrics but worse under others, leading to inconsistent rankings. Moreover, small specialists including ours sometimes exhibit deficiencies in fluency, as discussed in Sec. D.3, which existing metrics struggle to capture. As we examined, MLLM-as-judge yields relatively reliable results. In contrast, CLAIR showed higher variance across runs, limiting its consistency.
@@
D.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
We first provide a discussion, with examples, on how the inclusion of pseudo-initial captions leads to better supervision during training. We highlight the following points: (i) The model is trained to generate the ground-truth (GT) caption given both the image and the pseudo-initial caption as input: $I+\hat{c}\;\rightarrow\;\text{Our model}\;\rightarrow\;c$. (ii) If the pseudo-initial caption is generated following the strategies in Sec. F.1, then it is unlikely to include too many differing tokens from the GT caption.
Suppose we have “A woman in a room with a cat.” as the $c$ and $\hat{c}$ in Tab. 11.
pseudo-initial caption $\hat{c}$ different tokens $E={t\mid\hat{c}(t)\neq c(t)}$ in Sec. 5
@@ D.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
As an additional experiment, we fine-tune our captioner via Stage 2 as in Fig. 8 using pseudo-initial captions generated under the following four conditions: (Data 1) Initial captions generated by the captioner fine-tuned via Stage 1, (Data 2) Pseudo-initial captions that are significantly different from the GT caption, (Data 3) Pseudo-initial captions with minor modifications from the GT caption, which is our original strategy, and (Data 4) Pseudo-initial captions consisting of (i) one generated as in Data3, and (ii) the GT caption itself (e.g., GT: “A woman in a room with a cat”; pseudo-initials: “A boy in a room with a dog” and “A woman in a room with a cat”).
@@
Tab. 12 demonstrates that Data 1 provides reasonable gains within our framework, while Data 2 leads to the anticipated issue where the model tends to ignore the initial caption and generate a new one. Interestingly, Data 4 performs comparably to Data 3, and together with the results from Data 1, indicates that our framework is robust to moderate variation in pseudo-initial caption quality.
D.6 Further Research Questions
@@
    Do potential biases (e.g., gender, occupation, and age) exist in captioning models, and if so, how can we address them together with cultural contexts and deployment domains such as indoor, outdoor, and robotic environments? Could domain adaptation [64] with user feedback [62] serve as a solution?
@@
Figure 15: Additional results corresponding to the analyses in Fig. 5 and Fig. 6. It illustrates diffuse attention patterns (top) and the limited visual detail captured by CLIP features (bottom).
Appendix E Additional Experimental Results
@@
We present additional results from the analysis experiments conducted in Sec. 3, as shown in Fig. 15. These results support our earlier findings that single-pass captioners tend to adopt diffuse attention patterns, and that CLIP’s final-layer features may lack the granularity needed for fine-grained description.
@@
Additional results comparing our lightweight specialist with the large multimodal generalist (LLaVA-1.5-7B), as discussed in Sec. 3, are shown in Fig. 16. Despite its limited parameter size and a frozen vision encoder, our specialist demonstrates unexpectedly strong performance on captioning benchmarks, indicating that small models may still handle detailed captioning tasks reasonably.
@@
Additional results related to our refinement framework are presented in Fig. 17, building on the findings discussed in Sec. 6. Our refinement method yields caption-quality improvements in both single-sentence and detailed captioning.
Appendix F Extensive Details on The Dataset
F.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
We describe our strategy to generate the pseudo-initial caption as the following. The instruction prompt used for this strategy is shown in Fig. 19.
@@ F.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
  • Few-shot examples are provided like the examples in Tab. 13.
@@
Table 13: Few shot examples of GT captions and corresponding pseudo-initial captions.
We consistently generate three pseudo-initial captions per GT caption. Hence, the dataset can be summarized as Tab. 14, where ‘#pairs’ indicates the number of image-caption pairs.
@@
F.2 Datasets for Detailed Captioning
@@ F.4 GLaMM
Appendix G Prompt Templates
We share the prompt templates used when interacting with OpenAI’s GPT models throughout our study. First, the prompt used for summarizing long captions into shorter ones, introduced in Sec. F.3, is shown in Fig. 18. Second, the prompt used to generate pseudo initial captions for our refinement framework, as described in Sec. 5, is provided in Fig. 19. The examples generated using it are in Fig. 21. Lastly, the prompt used in the MLLM-as-judge evaluation setup is shown in Fig. 20. We note that we strictly follow the template proposed in the original MLLM-as-judge papers [8, 9].
@@ Appendix G Prompt Templates
Figure 21: Examples of pseudo initial captions generated using the prompt in Fig. 19, from MS COCO (top) and ShareGPT4V & DCI (bottom).
Appendix H Experimental Details
H.1 Pretraining and Finetuning
Our implementation is initialized from the LLaVA-1.5 repository [42]. To reduce reliance on LLMs, we replace LLaMA with the OPT series of language models [92]. The training process consists of three stages: pretraining, finetuning for caption generation (in Sec. 3), and finetuning for sharp-eyed refinement (in Sec. 5). The hyperparameters used for each stage are summarized in Tab. 15. We closely follow the original LLaVA training setup. As emphasized in Sec. D.1, our method differs from previous small-model approaches in that the visual features are directly used as self-attention inputs to the language model.
@@ H.1 Pretraining and Finetuning
Table 15: Hyperparameters used for model training. The settings to the left and right of the / correspond to those used in Sec. 4 and Sec. 3, respectively.
H.2 Experiment setting of large multimodal generalist
Tab. 2 compares specialist models with several generalist MLLMs, including InstructBLIP [15], Unified-IO-XL [46], Shikra [10], Qwen-VL [3], and LLaVA-1.5 [42]. These generalist models were reported to be trained on MS COCO images, but not on datasets such as ShareGPT4V, DCI, or GLaMM. Instead, they were instruction-tuned on large-scale datasets; for example, Qwen-VL and InstructBLIP were trained with approximately 1.5B and 130M instruction samples, respectively. For this reason, we categorize them as generalist models.
The generalist MLLMs were evaluated using publicly available checkpoints from their official GitHub repositories, without any additional fine-tuning. For the single-sentence captioning task, we used the prompt: “Provide a one-sentence caption for the provided image.” For the detailed captioning task, we used the prompt: “Describe the photo within 55 words.”. Additionally, we conduct an experiment on LLaVA-1.5 with the alternative instruction, as shown in Tab. 16. These results reveal two important points: (1) evaluation metrics such as BERTScore [93] penalize long-form outputs, and (2) longer generations tend to induce more hallucinations, consistent with prior observations [26].
@@ H.2 Experiment setting of large multimodal generalist
H.3 Attention map visualization
As part of our analysis in Sec. 4, we evaluate whether the model attends to the appropriate image regions when generating specific words in a caption. To this end, we adapt the visualization code provided by API [88], originally designed to point out attention maps between images and questions in VQA tasks. We modify the attention hooking module (https://github.com/yu-rp/apiprompting/blob/master/API/API_LLaVA/hook.py) to visualize attention between image regions and selected words within captions. This allows us to inspect which parts of the image the LM focuses on when producing certain tokens, such as the phrase ‘a white toilet’ in Fig. 5.
H.4 Image reconstruction
To investigate whether CLIP’s visual representations are coarse or ambiguous, we conducted an image reconstruction experiment using a Masked AutoEncoder (MAE) framework in Sec. 4. In this setup, a ViT encoder produces visual features, which the decoder then uses to reconstruct the image. We adopt the same visual encoder as used in LLaVA-1.5, ‘CLIP ViT-L/14-336’, and keep its parameters frozen. The decoder receives the visual embeddings without masking and predicts the corresponding RGB image. We utilize two types of visual inputs: (i) last-layer features (i.e., LF), and (ii) multi-level features (i.e., ML), where ML consists of outputs from layers 13, 18, and 23 of the encoder. The decoder is trained on 100K images from the MS COCO dataset. Further architectural details are available in the MAE repository, https://github.com/facebookresearch/mae/blob/main/models_mae.py. The hyperparameters for optimizing the decoder are summarized in the table below.
@@ H.4 Image reconstruction
H.5 Long Range Video Question Answering

### v2

@@ E.2 Limitations of Existing Multimodal Large Language Models
We examine the broad challenges faced by existing multimodal large language models (MLLMs), particularly in addressing the issue of visual blindness. Prior works such as Eyes Wide Shut [66] and Cambrian [65] have identified and attempted to mitigate this issue through the use of multiple vision encoders, including DINOv2 [54], SigLIPv2 [68], and CLIP [57]. However, as illustrated in Figure 12, even models with large parameter sizes continue to exhibit difficulty in generating consistent long-form captions in complex, multi-object scenes. We also conduct experiments using two recent capable MLLMs, LLaVA-Next [37] and LLaVA-OneVision [35]. Despite incorporating the sophisticated approach which is to split an image into grids and utilize all the features from each grid, both models continue to suffer from incorrect captioning results.
@@
These findings suggest that visual blindness is a persistent limitation across model scales and architectures. In this context, we believe that our Sharp-Eyed Refinement framework—which encourages attention to key regions via initial captions and leverages multi-level features from a single vision encoder—is a meaningful and efficient step toward addressing this challenge.
E.3 Limitations of lightweight captioners
@@
While we have demonstrated the potential of small models in captioning tasks, the use of lightweight LMs unavoidably introduces some limitations. In particular, we observe that the model occasionally suffers from issues such as repetitive phrasing, reduced fluency, limited OCR capability, and a lack of general world knowledge. Examples of these cases are provided in Figure 13. These limitations may stem from two primary factors: (i) the small number of parameters, which can restrict the model’s capacity for complex reasoning and language generation [30], and (ii) the limited scale and quality of training data, as our model was trained on approximately 500K image-caption pairs from ShareGPT-4V [11] which contains machine-generated captions. A natural direction for future work is to investigate how far the capabilities of small models can be scaled with access to larger and higher-quality training datasets. In addition to the results in Section E.2, we observe similar issues in larger models such as LLaVA-1.5, suggesting that these challenges remain unresolved [5, 66] and require deeper investigation.
E.4 Limitations of existing evaluation methods
@@
To ensure fair comparison across captioning models, we adopt seven evaluation metrics: BLEU@4 [56], METEOR [16], CIDEr [71], BERTScore [94], CAPTURE [18], CLAIR [8], and MLLM-as-judge [9] (i.e., GPT-based evaluation). For CLAIR and MLLM-as-judge, we randomly select 100 image samples and evaluate them using 10 different seeds to compute both mean and standard deviation. The values reported alongside the $\pm$ symbol indicate the standard deviation across seeds. Despite these efforts, we find that current evaluation methods do not always align well with human judgment. For instance, certain models perform better under specific metrics but worse under others, leading to inconsistent rankings. Moreover, small specialists including ours sometimes exhibit deficiencies in fluency, as discussed in Section E.3, which existing metrics struggle to capture. As we examined, MLLM-as-judge yields relatively reliable results. In contrast, CLAIR showed higher variance across runs, limiting its consistency.
@@
E.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
We first provide a discussion, with examples, on how the inclusion of pseudo-initial captions leads to better supervision during training. We highlight the following points: (i) The model is trained to generate the ground-truth (GT) caption given both the image and the pseudo-initial caption as input: $I+\hat{c}\;\rightarrow\;\text{Our model}\;\rightarrow\;c$. (ii) If the pseudo-initial caption is generated following the strategies in Section G.1, then it is unlikely to include too many differing tokens from the GT caption.
pseudo-initial caption $\hat{c}$ different tokens $E={t\mid\hat{c}(t)\neq c(t)}$ in Section 5
@@ E.5 Impact and Ablation of Pseudo-Initial Captions in Stage-2 Fine-tuning
Suppose we have “A woman in a room with a cat.” as the $c$ and $\hat{c}$ in Table 11.
@@
As an additional experiment, we fine-tune our captioner via Stage 2 as in Figure 7 using pseudo-initial captions generated under the following four conditions: (Data 1) Initial captions generated by the captioner fine-tuned via Stage 1, (Data 2) Pseudo-initial captions that are significantly different from the GT caption, (Data 3) Pseudo-initial captions with minor modifications from the GT caption, which is our original strategy, and (Data 4) Pseudo-initial captions consisting of (i) one generated as in Data3, and (ii) the GT caption itself (e.g., GT: “A woman in a room with a cat”; pseudo-initials: “A boy in a room with a dog” and “A woman in a room with a cat”).
@@
Table 12 demonstrates that Data 1 provides reasonable gains within our framework, while Data 2 leads to the anticipated issue where the model tends to ignore the initial caption and generate a new one. Interestingly, Data 4 performs comparably to Data 3, and together with the results from Data 1, indicates that our framework is robust to moderate variation in pseudo-initial caption quality.
E.6 Further Research Questions
@@
    Do potential biases (e.g., gender, occupation, and age) exist in captioning models, and if so, how can we address them together with cultural contexts and deployment domains such as indoor, outdoor, and robotic environments? Could domain adaptation with user feedback serve as a solution?
@@
Figure 14: Additional results corresponding to the analyses in Figure 5 and Figure 6. It illustrates diffuse attention patterns (top) and the limited visual detail captured by CLIP features (bottom).
Appendix F Additional Experimental Results
@@
We present additional results from the analysis experiments conducted in Section 3, as shown in Figure 14. These results support our earlier findings that single-pass captioners tend to adopt diffuse attention patterns, and that CLIP’s final-layer features may lack the granularity needed for fine-grained description.
@@
Additional results comparing our lightweight specialist with the large multimodal generalist (LLaVA-1.5-7B), as discussed in Section 3, are shown in Figure 15. Despite its limited parameter size and a frozen vision encoder, our specialist demonstrates unexpectedly strong performance on captioning benchmarks, indicating that small models may still handle detailed captioning tasks reasonably.
@@
Additional results related to our refinement framework are presented in Figure 16, building on the findings discussed in Section 6. Our refinement method yields caption-quality improvements in both single-sentence and detailed captioning.
Appendix G Extensive Details on The Dataset
G.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
We describe our strategy to generate the pseudo-initial caption as the following. The instruction prompt used for this strategy is shown in Figure 18.
@@ G.1 How to generate the pseudo-initial caption $\hat{c}_{k}$?
  • Few-shot examples are provided like the examples in Table 13.
@@
Table 13: Few-shot examples of GT captions and corresponding pseudo-initial captions.
We consistently generate three pseudo-initial captions per GT caption. Hence, the dataset can be summarized as Table 14, where ‘#pairs’ indicates the number of image-caption pairs.
@@
G.2 Datasets for Detailed Captioning
@@ G.4 GLaMM
Appendix H Prompt Templates
We share the prompt templates used when interacting with OpenAI’s GPT models throughout our study. First, the prompt used for summarizing long captions into shorter ones, introduced in Section G.3, is shown in Figure 17. Second, the prompt used to generate pseudo initial captions for our refinement framework, as described in Section 5, is provided in Figure 18. The examples generated using it are in Figure 20. Lastly, the prompt used in the MLLM-as-judge evaluation setup is shown in Figure 19. We note that we strictly follow the template proposed in the original MLLM-as-judge papers [8, 9].
@@ Appendix H Prompt Templates
Figure 20: Examples of pseudo initial captions generated using the prompt in Figure 18, from MS COCO (top) and ShareGPT4V & DCI (bottom).
Appendix I Experimental Details
I.1 Pretraining and Finetuning
Our implementation is initialized from the LLaVA-1.5 repository [43]. To reduce reliance on LLMs, we replace LLaMA with the OPT series of language models [93]. The training process consists of three stages: pretraining, finetuning for caption generation (in Section 3), and finetuning for sharp-eyed refinement (in Section 5). The hyperparameters used for each stage are summarized in Table 15. We closely follow the original LLaVA training setup. As emphasized in Section E.1, our method differs from previous small-model approaches in that the visual features are directly used as self-attention inputs to the language model.
@@ I.1 Pretraining and Finetuning
Table 15: Hyperparameters used for model training. The settings to the left and right of the / correspond to those used in Section 4 and Section 3, respectively.
I.2 Experiment setting of large multimodal generalist
Table 2 compares specialist models with several generalist MLLMs, including InstructBLIP [15], Unified-IO-XL [47], Shikra [10], Qwen-VL [3], and LLaVA-1.5 [43]. These generalist models were reported to be trained on MS COCO images, but not on datasets such as ShareGPT4V, DCI, or GLaMM. Instead, they were instruction-tuned on large-scale datasets; for example, Qwen-VL and InstructBLIP were trained with approximately 1.5B and 130M instruction samples, respectively. For this reason, we categorize them as generalist models.
The generalist MLLMs were evaluated using publicly available checkpoints from their official GitHub repositories, without any additional fine-tuning. For the single-sentence captioning task, we used the prompt: “Provide a one-sentence caption for the provided image.” For the detailed captioning task, we used the prompt: “Describe the photo within 55 words.”. Additionally, we conduct an experiment on LLaVA-1.5 with the alternative instruction, as shown in Table 16. These results reveal two important points: (1) evaluation metrics such as BERTScore [94] penalize long-form outputs, and (2) longer generations tend to induce more hallucinations, consistent with prior observations [26].
@@ I.2 Experiment setting of large multimodal generalist
I.3 Attention map visualization
As part of our analysis in Section 4, we evaluate whether the model attends to the appropriate image regions when generating specific words in a caption. To this end, we adapt the visualization code provided by API [89], originally designed to point out attention maps between images and questions in VQA tasks. We modify the attention hooking module (https://github.com/yu-rp/apiprompting/blob/master/API/API_LLaVA/hook.py) to visualize attention between image regions and selected words within captions. This allows us to inspect which parts of the image the LM focuses on when producing certain tokens, such as the phrase ‘a white toilet’ in Figure 5.
I.4 Image reconstruction
To investigate whether CLIP’s visual representations are coarse or ambiguous, we conducted an image reconstruction experiment using a Masked AutoEncoder (MAE) framework in Section 4. In this setup, a ViT encoder produces visual features, which the decoder then uses to reconstruct the image. We adopt the same visual encoder as used in LLaVA-1.5, ‘CLIP ViT-L/14-336’, and keep its parameters frozen. The decoder receives the visual embeddings without masking and predicts the corresponding RGB image. We utilize two types of visual inputs: (i) last-layer features (i.e., LF), and (ii) multi-level features (i.e., ML), where ML consists of outputs from layers 13, 18, and 23 of the encoder. The decoder is trained on 100K images from the MS COCO dataset. Further architectural details are available in the MAE repository, https://github.com/facebookresearch/mae/blob/main/models_mae.py. The hyperparameters for optimizing the decoder are summarized in the table below.
@@ I.4 Image reconstruction
I.5 Long Range Video Question Answering